OpenStudy (anonymous):

PB has endpoints at -2 and 12. what is the coordinates of its midpoint...2 points!!!! Need serious help!

OpenStudy (anonymous):

This is hard!

OpenStudy (anonymous):

(0,-2) (0,12)

OpenStudy (phi):

if you plotted the two end points, then count up from one to the other, the mid point will be half way

OpenStudy (anonymous):

So what is it?

OpenStudy (anonymous):

i solved it already right?

OpenStudy (phi):

are you asking what number is half way between -2 and 12 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(0,2)(12,0)

OpenStudy (phi):

write down on paper -2 -1 0 1 ... up to 12 then count half way

OpenStudy (anonymous):

not (0,12)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

therfore (0+12)/2 =6

OpenStudy (anonymous):

and (-2+0)/2 = -1

OpenStudy (anonymous):

co-ordinates =(6,-1)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

k